OpenStudy (anonymous):

i have to simplify the equation: 1/8+2/t=17/8t

OpenStudy (anonymous):

t=-38/9

OpenStudy (anonymous):

can you show me the work?

OpenStudy (anonymous):

1/8=17/8t-2/t 1/8=(17-16)/8t 1/8=1/8t t=1

OpenStudy (anonymous):

srri ans 1 n nt -38/9

OpenStudy (anonymous):

right, thanks!